OpenStudy (anonymous):

find the solution of this system of equation. I need the solution for x and y -6x+2y=38 and 6x-8y=-44

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[(-6x+2y)+(6x-8y)=38+(-44)\] \[-6y=-6\] \[y=1\]

OpenStudy (anonymous):

sub in y to find x

OpenStudy (anonymous):

so what is x?

OpenStudy (anonymous):

6x-8(1)=-44 6x=-36

OpenStudy (anonymous):

x=6